You are provided with the following information for Concord Inc. for the month ended June 30, 2019. Concord uses the periodic system for inventory.

Date

Description

Quantity

Unit Cost or
Selling Price

June 1 Beginning inventory 43 $41
June 4 Purchase 137 45
June 10 Sale 115 70
June 11 Sale return 17 70
June 18 Purchase 56 47
June 18 Purchase return 8 47
June 25 Sale 67 75
June 28 Purchase 35 51

(a1)Calculate weighted-average cost per unit. (Round answer to 2 decimal places, e.g. 5.25.)

Weighted-average cost per unit

$

(a2)

Calculate ending inventory, cost of goods sold, gross profit under each of the following methods. (1) LIFO. (2) FIFO. (3) Average-cost. (Round average-cost method answers to 2 decimal places, e.g. 1,250.25 and other answers to 0 decimal places, e.g. 1,250. Use weighted-average unit cost rounded to 2 decimal places for computations.)

LIFO

FIFO

AVERAGE-COST

The ending inventory

$

$

$

The cost of goods sold

$

$

$

Gross profit

$

$

$
Answer #1

(A1) weighted average cost per unit

Date Description Quantity unit cost total cost
June 1 beginning inventory 43 41 1,763
June 4 purchase 137 45 6,165
June 18 purchase 56 47 2,632
June 18 Purchase return (8) 47 (376)
June 28 purchase 35 51 1,785
Total 263 11,969

Weighted average cost per unit= 45.51 per unit

A2) 1 ) lifo method

COGS Ending inventory

Date description quantity quantity per unit cost total cost quantity per unit cost total cost
June 1 beginning inventory 43 $ 41 $ 1,763
June 4 Purcahse 137 43 41 1,763
137 45 6,165
June 10 sales 115 115 45 5,175 43 41 1,763
22 45 990
June 11 Sales return 17 (17) 45 (765) 43 41 1,763
39 45 1,755
June 18 purchase 56 43 41 1,763
39 45 1,755
56 47 2,632
June 18 Purcahse return 8 43 41 1,763
39 45 1,755
48 47 2,256
June 25 sales 67 48 47 2,256 43 41 1763
19 45 855 20 45 900
June 28 purchase 35 43 41 1763
20 45 900
35 51 1785
Total 165 7,521 98 4,448

Total sales

Date description quantity sale price total sales
June 10 sales 115 70 8,050
June 11 sales return (17) 70 (1,190)
June 25 sales 67 75 5,025
Total 165 11,885

Ending inventory= 4,448

Cost of goods sold=7,521

Gross profit = sales - COGS = 11,885-7,521= 4,364

A2)2 FIFO Method

Date description quantity quantity per unit cost total cost quantity per unit cost total cost
June 1 beginning inventory 43 41 1,763
June 4 purchase 137 43 41 1,763
137 45 6,165
June 10 sale 115 43 41 1763
72 45 3,240 65 45 2,925
June 11 sales return 17 (17) 45 (765) 82 45 3,690
June 18 purchase 56 82 45 3,690
56 47 2,632
June 18 purchase return (8) 82 45 3,690
48 47 2,256
June 25 Sale 67 67 45 3015 15 45 675
48 47 2,256
June 28 purchase 35 15 45 675
48 47 2,256
35 51 1,785
Total 165 7,253 98 4,716

Ending inventory= 4,716

Cost of goods sold=7,253

Gross profit= 11,885 - 7,253= 4,632

Average cost

Ending inventory = 98 units × 45.51= 4,459.98

Cost of goods sold= 165 units × 45.51= 7,509.15

Gross profit= 11,885-7509.15 = 4,375.85